Through which force does the photon interact with
Answer #1

the photon is a quantum of light, in word, is the particle form of light, therefore is interacting with the other particles through the electromagnetic forces, 'cause the light obeys the Maxwell's equations.

Maxwell's equations are (in vacuum):

i)\;\;\nabla \cdot \vec E = 0

ii)\;\;\nabla \cdot \vec B = 0

iv)\;\;\nabla \times \vec E = -\frac{\partial B}{\partial t}

v)\;\;\nabla \times \vec B = \mu_o\epsilon_o\frac{\partial E}{\partial t}

thus:

\nabla^2 E = \frac{1}{c^2}\frac{\partial^2E}{\partial t^2}

\nabla^2 B = \frac{1}{c^2}\frac{\partial^2B}{\partial t^2}